Circular No. 8773

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2006my–2006ne S. Nakano, Sumoto, Japan, reports the discovery by K. Itagaki (Teppo-

cho, Yamagata, Japan, 0.60-m f/5.7 reflector) of an apparent supernova (mag 15.3) on unfiltered CCD exposures taken around Nov. 8.82 and 9.81 UT; 2006my is located at α = 12^{h}43^{m}40^{s}.74, δ = +16^{o}23′14′′.1 (equinox 2000.0), which is 27′′ west and 22′′.5 south of the center of NGC 4651. Nothing is visible at this location on the Digitized Sky Survey or on Itagaki’s many past CCD frames, including those taken on Feb. 24 (limiting mag 19.5) and June 28 (limiting mag 19.0). SN 2006my is a type-II supernova (details on CBET 737), well past maximum light.

J. Frieman and the Sloan Digital Sky Survey II collaboration report the discovery of six additional faint type-Ia supernovae, designated 2006mz– 2006ne (g-magnitude range 21.1–22.8), with details given on CBET 735.

R/2006 S 5 E. Roussos, G. H. Jones, and N. Krupp, Max-Planck-Institut für Son-

nensystemforschung, Katlenburg-Lindau; and S. M. Krimigis, D. Mitchell, and C. Paranicas, Applied Physics Laboratory, Johns Hopkins University, report the probable detection of a ring arc (designated R/2006 S 5) around the orbital distance of the small satellite Saturn XXXII (Methone; IAUC 8339). The feature’s existence is indicated by two energetic electron depletions detected by the LEMMS sensor of the Cassini Magnetospheric Imaging Instrument (MIMI) on 2006 Sept. 9, at equivalent equatorial distances of 197700 and 194700 km — some radial shifting in the signature being possible due to magnetospheric effects. Methone’s distance is 194000 km. The depletions’ widths at the equator are around 1000 km — inconsistent with absorption by the ≈ 3-km-wide Methone. The feature’s nature as an arc is suggested by the spacecraft’s orbital geometry with respect to Methone and the absence of such absorption signatures at other crossings of the satellite’s orbit. The arc’s longitudinal extent appears possibly to include the position of Methone itself.

COMETS 179P/2006 U2, 180P/2006 U3, AND 181P/2006 U4 Three newly recovered short-period comets have been assigned perma-

nent numberings (cf. MPC 57901): 179P/Jedicke (= P/2006 U2; cf. IAUC 8764); 180P/NEAT (= P/2006 U3; cf. IAUC 8765); 181P/Shoemaker-Levy (= P/2006 U4; cf. IAUC 8767).

2006 November 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green